# On K-peak solutions for the Yamabe equation on product manifolds

### 09 Mar 2026 v1

@@ Abstract.
Let $(M^{n},g)$ and $(X^{m},h)$ be closed manifolds $m,n>2$, such that $(X,h)$ has constant positive scalar curvature. We consider the one parameter family of products $(M\times X,g+\epsilon^{2}h)$, $\epsilon>0$. Let $\xi_{0}\in M$ be a stable critical point of a function $\Phi:M\rightarrow{\mathbb{R}}$, eq. (9), that depends on the scalar curvature, the dimensions $m,n$, the norm of the Ricci curvature and the norm of the curvature tensor of $g$. We prove that, if either the scalar curvature of $g$ is constant or a certain dimensional constant $\beta=0$, then for each $K\in{\mathbb{N}}$, there is some $\epsilon_{0}>0$ such that for every $\epsilon\in(0,\epsilon_{0})$ the subcritical Yamabe equation $\epsilon^{2}\Delta u+(1+{\bf{c}}\epsilon^{2}s_{g})u=u^{q}$ on $(M\times X,g+\epsilon^{2}h)$ has a $K$-peak positive solution. Where, ${\bf{c}}=\frac{N-2}{4(N-1)}$, $N=n+m$, $q=\frac{N+2}{N-2}$ and $s_{g}$ the scalar curvature of $(M,g)$. This covers some remaining cases of previous results [18] and gives examples of multiplicity of positive solutions for the Yamabe equation.
@@ 1. Introduction
where $a_{n}=\frac{4(n-1)}{n-2}$ and $p_{n}=\frac{2n}{n-2}$. It is a classical result, solved in various steps by H. Yamabe [22], N.S. Trudinger [21], T. Aubin [2], and R. Schoen [20], that such a postive, smooth, function $u$ always exists (see [15], [1], for a detailed discussion of the Yamabe problem). It has been of interest to ask whether other metrics of constant scalar curvature exist in the same conformal class, this is, whether problem (1) has a unique solution. Unicity in this sense is known to be true for some special cases, for example, when the original metric has non-positive scalar curvature, by the maximum principle [2], or when the original metric is an Einstein metric, not isometric to the sphere with the round metric, by the work of M. Obata [12]. On the other hand, when the original metric has positive scalar curvature, many examples of different metrics with constant scalar curvature and unit volume have been found in the same conformal class, see for instance [3], [6], [9], [13], [14], [16], [17], [19].
@@
where ${\bf{c}}=c_{N}=a_{N}^{-1}$. To solve (3) , we follow the techniques of the Lyapunov-Schmidt reduction, which have been applied by many authors for these type of equations. We mention in particular the work of Dancer, Micheletti and Pistoia [4], see also [18] and the work of Deng, Khemiri and Mahmoudi [7].
@@
Note that $W_{\epsilon,\xi}(x)$ is an approximate solution to (3), which concentrates around $\xi$ as $\epsilon\rightarrow 0$, by the exponential decay of $U$, eventually getting closer to an exact solution. Adding $K$ of these peaks $W_{\epsilon,\xi}(x)$, centered at different $K$ strategic points $\xi_{1},\xi_{2},\dots,\xi_{K}\in M$, we get an approximate solution of (3) which has $K$-peaks: $\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}(x)$.
Then, by adding some perturbations of these approximate solutions we will get an exact solution to equation (3).
@@
If $\beta\neq 0$ and $s_{g}$ has isolated, stable, critical points, multipeak solutions like the ones described above were studied in [18]. In this case, the solutions concentrate around isolated critical points of the scalar curvature $s_{g}$. In the present article we work on the remaining cases: if either $\beta=0$ or $s_{g}$ is constant. In particular, we show that, in these cases, in order to construct these multiple peak solutions, one needs to look instead for an isolated, stable, critical point of the functional $\Phi:M\rightarrow{\mathbb{R}}^{n}$, given by
@@
where $s_{g}$, $Ric_{\xi}$ and $R_{\xi}$ are the scalar curvature, Ricci curvature and tensor of curvature of $(M,g)$, respectively. Meanwhile $c_{1},c_{6},c_{7},c_{8}$ and $c_{9}$ are constants that depend on the dimensions $m,n$, and are given in equations (71) through (79). Note that, for example, if $(M,g)$ is an Einstein manifold that does not have constant sectional curvature, the solutions concentrate around $C^{1}$ stable critical points of $R_{\xi}||$. We also remark that at least the constant $c_{1}=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz\neq 0$ for any values of $n,m>2$.
@@
We follow the ideas in [8], of expanding an energy functional for problem (3) up to order 4, which gives place to the functional $\Phi(\xi)$. We also follow standard Lyapunov-Schmidt reduction techinques for subcritical elliptic PDEs on manifolds, developed for example in [4] and [11]. We also follow [18], where these techniques were applied to the subcritical Yamabe equation.
@@ 2. Notation and background
We now look for a second order improvement of the approximation solution $W_{\epsilon,\xi}$ of the form
@@
where $V_{\epsilon,\xi}$ will be a function $V_{\epsilon,\xi}:M\rightarrow{\mathbb{R}}$, such that the second order terms of $\epsilon$ vanish in
@@
As computed in [8], section 3.1, $\frac{1}{3}R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U=\frac{1}{3}R_{kl}z_{k}\ \partial_{l}U$, hence eq. (22) is equivalent to
@@
Note that (25) is always solvable since the right-hand side is orthogonal to the Kernel of $L_{0}$ (which is spanned by $\frac{\partial U}{\partial x_{i}}$, $i=1,2,\dots,n$). As developed in [8], section 3.1, one can look for the solution $V_{1}$ to (25) of the form
(27) $V_{1}=\frac{1}{3}\left(\sum_{k\neq l}^{n}R_{kl}(\xi)z_{k}z_{l}\psi_{1}(|z|)+\frac{1}{n}s_{g}(\xi)\psi_{2}(|z|)\right)$
where $\psi_{1}$ and $\psi_{2}$ are radial functions; since equation (25) is equivalent to
@@
and $\frac{U^{\prime}}{|z|}$ is radial. Moreover, by the form of the operator $L_{0}$ (eq. (23)) and of the right hand side of (25), $V_{1}$ and its derivatives are exponentially decaying at infinity.
@@
satisfies (26). Of course, this implies that $V=V_{1}+V_{2}$ and its derivatives are exponentially decaying at infinity.
@@
Let $\xi_{0}\in M$ an isolated local stable critical point of the functional $\Phi$ (eq. (9)). Let $K\geq 1$ be an integer. Given $\rho>0$, $\epsilon>0$, let
@@
since, otherwise, if for some $a>0$,
(37) $\lim_{\epsilon\rightarrow 0}e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4},$
then, for some small and positive values of $\epsilon$ ($\epsilon\rightarrow 0$),
(38) $\lim_{\epsilon\rightarrow 0}e^{-\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4}e^{\delta\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}},$
@@
which gives a contradiction as $\epsilon\rightarrow 0$, since one has polynomial growth and the other exponential growth. In a similar way, note that, for $\delta>0$ small, also
@@
since, otherwise, if for some $b>0$
(41) $\lim_{\epsilon\rightarrow 0}e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4},$
for some small and positive values of $\epsilon$, as $\epsilon\rightarrow 0$, then for those values
(42) $\lim_{\epsilon\rightarrow 0}e^{-\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4}e^{(\frac{1}{2}-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}},$
@@
which gives a contradiction, for small $\delta$, as $\epsilon\rightarrow 0$, since the left hand side has polynomial growth and the other side exponential growth. Equations (36) and (40) imply that $e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ and $e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ are $o(\epsilon^{4})$ for ${\xi_{\epsilon}}_{i}$, ${\xi_{\epsilon}}_{j}\in\bar{\xi}_{\epsilon}$, with $\bar{\xi}_{\epsilon}\in D^{k}_{\epsilon,\rho}$.
@@
Let $S_{\epsilon}=\nabla J_{\epsilon}:H_{\epsilon}\rightarrow H_{\epsilon}$. Note that a solution to (15) is a critical point to $J_{\epsilon}$ (eq. (44)) and hence a positive function.
To solve problem (3) we must solve $S_{\epsilon}(u)=0$. Or in an equivalent way, we must solve the system,
(45) $\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi)\right\}=0.$
(46) $\Pi_{\epsilon,\xi}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi)\right\}=0.$
@@
The proof of this proposition is presented in section 5.
For $\bar{\xi}\in M^{K}$, recall that $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})$, where $W_{\epsilon,\xi_{i}}$ and $V_{\epsilon,\xi_{i}}$ are defined in (7) and (21), respectively. We define the functional $\bar{J}_{\epsilon}:D_{\epsilon,\rho}^{K}\subset M^{K}\rightarrow{\mathbb{R}}$ by
@@
where $c_{1},c_{6},c_{7}$ and $c_{8}$ are constants that depend on the dimensions $m,n$, and are given in equations (71) through (79).
Note that, in particular, from (53): $|b_{ij}|=1$. The proof of this proposition is presented in section 4.
@@
If $\bar{\xi}$ is a critical point of $\bar{J}_{\epsilon}$, then the function $Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}$ is a solution to equation (46), or, equivalently, to problem (3).
We present the proof of this proposition to section 6. We now prove Theorem 1.1.
@@
We first note that by eq. (49), for $\bar{\xi}_{\epsilon}$,
@@
for $\epsilon$ small enough, since $U$ is radially symetric and of exponential decay. This implies $\bar{\eta}_{\epsilon}\in D^{K}_{\epsilon,\rho}$ holds, if $\epsilon$ is small enough.
By (57) and (49)
@@
(59) $\Phi(\eta_{i})=\Phi(\xi_{0})+\frac{1}{2}\Phi^{\prime\prime}(\xi_{0})d(\xi_{0},\eta_{i})^{2}+o(\epsilon)=\Phi(\xi_{0})+\Phi^{\prime\prime}(\xi_{0})\epsilon|e_{i}|^{2}+o(\epsilon)$
@@
where
$\gamma:=\min\left\{\int_{{\mathbb{R}}^{n}}U^{p-1}(z)e^{\langle b_{ij},z\rangle}dz|b_{ij}\in{\mathbb{R}}^{n},|b_{ij}|=1\right\}>0.$

### 19 Mar 2026 v2

@@ Abstract.
Let $(M^{n},g)$ and $(X^{m},h)$ be closed manifolds $m,n>2$, such that $(X,h)$ has constant positive scalar curvature. We consider the one parameter family of products $(M\times X,g+\epsilon^{2}h)$, $\epsilon>0$. Let $\xi_{0}\in M$ be a stable, isolated, critical point of a function $\Phi:M\rightarrow{\mathbb{R}}$, eq. (9), that depends on the scalar curvature, the dimensions $m,n$, the norm of the Ricci curvature and the norm of the curvature tensor of $g$. We prove that, if either the scalar curvature of $g$ is constant or a certain dimensional constant $\beta=0$, then for each $K\in{\mathbb{N}}$, there is some $\epsilon_{0}>0$ such that for every $\epsilon\in(0,\epsilon_{0})$ the subcritical Yamabe equation $\epsilon^{2}\Delta u+(1+{\bf{c}}\epsilon^{2}s_{g})u=u^{q}$ on $(M\times X,g+\epsilon^{2}h)$ has a positive $K$-peak solution, which concentrates around $\xi_{0}$. Where, ${\bf{c}}=\frac{N-2}{4(N-1)}$, $N=n+m$, $q=\frac{N+2}{N-2}$ and $s_{g}$ the scalar curvature of $(M,g)$. This covers some remaining cases of previous results [19] and gives examples of multiplicity of positive solutions for the Yamabe equation.
@@ 1. Introduction
where $a_{n}=\frac{4(n-1)}{n-2}$ and $p_{n}=\frac{2n}{n-2}$. It is a classical result, solved in various steps by H. Yamabe [23], N.S. Trudinger [22], T. Aubin [2], and R. Schoen [21], that such a postive, smooth, function $u$ always exists (see [16], [1], for a detailed discussion of the Yamabe problem). It has been of interest to ask whether other metrics of constant scalar curvature exist in the same conformal class, this is, whether problem (1) has a unique solution. Uniqueness in this sense is known to be true for some special cases, for example, when the original metric has non-positive scalar curvature, by the maximum principle [2], or when the original metric is an Einstein metric, not isometric to the sphere with the round metric, by the work of M. Obata [13]. On the other hand, when the original metric has positive scalar curvature, many examples of different metrics with constant scalar curvature and unit volume have been found in the same conformal class, see for instance [3], [6], [10], [14], [15], [17], [18], [20].
@@
where ${\bf{c}}=c_{N}=a_{N}^{-1}$. To solve (3), we follow the techniques of the Lyapunov-Schmidt reduction, which have been applied by many authors for these type of equations. We refer the reader, in particular, to the work of Dancer, Micheletti and Pistoia [4], of Rey and Ruiz [19] and of Deng, Khemiri and Mahmoudi [7].
@@
Note that $W_{\epsilon,\xi}(x)$ is an approximate solution to (3), which concentrates around $\xi$ as $\epsilon\rightarrow 0$, by the exponential decay of $U$, eventually getting closer to an exact solution. Adding $K$ of these peaks $W_{\epsilon,\xi}(x)$, centered at different $K$ strategic points $\xi_{1},\xi_{2},\dots,\xi_{K}\in M$, we get an approximate solution of (3) which has $K$-peaks: $\sum_{i=1}^{K}W_{\epsilon,\xi_{i}}(x)$. Then, by adding some perturbations of these approximate solutions we will get an exact solution to equation (3).
@@
If $\beta\neq 0$ and $s_{g}$ has isolated, stable, critical points, multipeak solutions like the ones described above were studied in [19]. In this case, the solutions concentrate around isolated critical points of the scalar curvature $s_{g}$. In the present article we work on some of the remaining cases. In particular, we show that, if either $\beta=0$ or $s_{g}$ is constant, $K$-peak solutions concentrate around an isolated, stable, critical point of the functional $\Phi:M\rightarrow{\mathbb{R}}^{n}$, given by
@@
where $s_{g}$, $Ric_{\xi}$ and $R_{\xi}$ are the scalar curvature, Ricci curvature and tensor of curvature of $(M,g)$, respectively. Meanwhile $c_{1},c_{6},c_{7},c_{8}$ and $c_{9}$ are constants that depend on the dimensions $m,n$, and are given in equations (69) through (74). Note that, for example, if $(M,g)$ is an Einstein manifold, then critical points of $R_{\xi}||^{2}$ would be critical points of $\Phi(\xi)$. We also remark that at least the constant $c_{1}=\frac{1}{6}\int_{{\mathbb{R}}^{n}}\left(\frac{U^{\prime}(|z|)}{|z|}\right)^{2}z_{1}^{4}dz\neq 0$ for any values of $n,m>2$.
@@
Our approach follows some ideas in [8], like expanding the energy functional to order 4, which gives place to the functional $\Phi(\xi)$. We then use Lyapunov-Schmidt reduction techinques for subcritical elliptic PDEs on manifolds, like the ones applied, for example, in [12], [7]; in particular [4] for multipeak solutions and [19] for the subcritical Yamabe equation.
In Section 2 we introduce some notation and background, as well as a way to improve the aproximate solution to higher order. In section 3 we prove Theorem 1.1, assuming propositions 3.1, 3.2 and 3.3. Section 4 is dedicated to compute the expansion of the energy functional, which is the content of proposition 3.2. In section 5 we prove proposition 3.1, regarding the finite dimensional reduction. Finally, in section 6, we provide some technical estimates and the proof of proposition 3.3.
@@ 2. Approximate solutions
The following will be useful. They are computed, for example, in [9].
@@
We will use a second order improvement of the approximation solution $W_{\epsilon,\xi}$ of the form
@@
where $V_{\epsilon,\xi}$ will be a function $V_{\epsilon,\xi}:M\rightarrow{\mathbb{R}}$, such that the second order terms of $\epsilon$ vanish for
@@
Since $U$ is radial,
$\partial_{i}U=z_{i}\frac{U^{\prime}}{|z|}\textnormal{\ \ \ and \ \ \ }\partial_{ij}^{2}U=\delta_{ij}\frac{U^{\prime\prime}}{|z|^{2}}+z_{i}z_{j}\frac{U^{\prime}}{|z|}-z_{i}z_{j}\frac{U^{\prime}}{|z|^{3}}$
then
$R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U=R_{kiil}z_{k}z_{l}\frac{U^{\prime}}{|z|}+R_{kijl}z_{k}z_{l}z_{i}z_{j}\frac{U^{\prime\prime}}{|z|^{2}}-R_{kijl}z_{k}z_{l}z_{i}z_{j}\frac{U^{\prime}}{|z|^{3}}$
with the second and third terms being zero, by the antisymmetry of the curvature tensor. Thus, we have,
$\frac{1}{3}R_{kijl}z_{k}z_{l}\partial_{ij}^{2}U+\frac{2}{3}R_{kssj}z_{k}\partial_{j}U=\frac{1}{3}R_{kl}z_{k}\ \partial_{l}U$
this implies that eq. (22) is equivalent to
@@
Note that (25) is always solvable since the right-hand side is orthogonal to the Kernel of $L_{0}$ (which is spanned by $\partial_{i}U$, $i=1,2,\dots,n$). Moreover, since $U$ is radial, then $\partial_{l}U=\frac{U^{\prime}}{|z|}z_{l}$, hence eq. (25) is equivalent to
@@
By direct computation, for a radial function $\varphi=\varphi(|z|)$ on ${\mathbb{R}}^{n}$,
$L(\varphi z_{k}z_{l})=\left(\varphi^{\prime\prime}+(N+3)\frac{\varphi^{\prime}}{|z|}+\varphi-(p-1)U^{p-2}\varphi\right)z_{k}z_{l}$
Then, one can look for a radial function $\psi=\psi(|z|)$, such that
(28) $L(\psi z_{k}z_{l})=\frac{U^{\prime}}{|z|}z_{k}z_{l}$
since $\frac{U^{\prime}}{|z|}$ is radial. In this way, the solution to (27) will be given by
(29) $V_{1}=\frac{1}{3}R_{kl}\ \psi\ z_{k}z_{l}$
Moreover, by the form of the operator $L_{0}$ (eq. (23)) and of the right hand side of (27), $V_{1}$ and its derivatives are exponentially decaying at infinity.
@@
satisfies (26). Of course, this implies that $V=V_{1}+V_{2}$ satisfies (24) and, moreover, that $V$ and its derivatives are exponentially decaying at infinity.
@@
Let $\xi_{0}\in M$ an isolated, stable, $C^{1}$ critical point of the functional $\Phi$ (eq. (9)). Let $K\geq 1$ be an integer. Given $\rho>0$, $\epsilon>0$, let
@@
since, otherwise, if for some $a>0$, $\lim_{\epsilon\rightarrow 0}e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4}$. Then, for some small and positive values of $\epsilon$ ($\epsilon\rightarrow 0$),
$\lim_{\epsilon\rightarrow 0}e^{-\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>a\epsilon^{4}e^{\delta\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}.$
@@
which gives a contradiction as $\epsilon\rightarrow 0$, since one side has polynomial growth and the other one exponential growth. In a similar way, note that, for $\delta>0$ small, also
@@
since, otherwise, if for some $b>0$, $\lim_{\epsilon\rightarrow 0}e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4}$, for some small and positive values of $\epsilon$, as $\epsilon\rightarrow 0$. And then, for those values,
(39) $\lim_{\epsilon\rightarrow 0}e^{-\frac{1}{2}\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}>b\epsilon^{4}e^{(\frac{1}{2}-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}.$
@@
which gives a contradiction, for small $\delta$, as $\epsilon\rightarrow 0$, since one side has polynomial growth and the other, exponential growth. Equations (36) and (38) imply that $e^{-(1+\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ and $e^{-(1-\delta)\frac{d_{g}({\xi_{\epsilon}}_{i},{\xi_{\epsilon}}_{j})}{\epsilon}}$ are $o(\epsilon^{4})$ for ${\xi_{\epsilon}}_{i}$, ${\xi_{\epsilon}}_{j}\in\bar{\xi}_{\epsilon}$, $i\neq j$, with $\bar{\xi}_{\epsilon}\in D^{k}_{\epsilon,\rho}$.
We will look for a solution to (3) of the form
(41) $u_{\epsilon}=Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}$
where $Y_{\epsilon,\bar{\xi}}$ is defined in (33), $\bar{\xi}_{\epsilon}\in D^{k}_{\epsilon,\rho}$ and the rest term $\phi_{\epsilon,\bar{\xi}}\in K^{\perp}_{\epsilon,\bar{\xi}}$.
@@
Let $S_{\epsilon}=\nabla J_{\epsilon}:H_{\epsilon}\rightarrow H_{\epsilon}$. To solve problem (3) we must solve $S_{\epsilon}(u_{\epsilon})=0$. Or in an equivalent way, we must solve the system,
(43) $\Pi_{\epsilon,\xi}^{\perp}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})\right\}=0.$
(44) $\Pi_{\epsilon,\xi}\left\{S_{\epsilon}(Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}})\right\}=0.$
@@
We will use the following to prove Theorem 1.1. The proof of this proposition is postponed to section 5.
@@
For $\bar{\xi}\in M^{K}$, recall that $Y_{\epsilon,\bar{\xi}}=\sum_{i=1}^{K}(W_{\epsilon,\xi_{i}}+\epsilon^{2}V_{\epsilon,\xi_{i}})$, $W_{\epsilon,\xi_{i}}$ and $V_{\epsilon,\xi_{i}}$ being defined in (7) and (21), respectively. We define the functional $\bar{J}_{\epsilon}:D_{\epsilon,\rho}^{K}\subset M^{K}\rightarrow{\mathbb{R}}$, by
@@
The following gives the expansion of the energy functional. Its proof is presented in section 4.
@@
where $c_{1},c_{6},c_{7},c_{8}$ and $c_{9}$ are constants that depend on the dimensions $m,n$, and are given in equations (69) through (74).
Note that, in particular, from (51): $|b_{ij}|=1$. Note next that critical points of $\bar{J}_{\epsilon}$ are solutions to problem (44), by the following.
@@
If $\bar{\xi}$ is an isolated, stable, $C^{1}$ critical point of $\bar{J}_{\epsilon}$, then the function $Y_{\epsilon,\bar{\xi}}+\phi_{\epsilon,\bar{\xi}}$ is a solution to equation (44), or, equivalently, to problem (3).
We present the proof of this proposition in section 6. We now prove Theorem 1.1.
@@
We first note that by Proposition 3.2, eq. (47), for $\bar{\xi}_{\epsilon}$:
@@
for $\epsilon$ small enough, since $U$ is radially symetric and of exponential decay. This implies $\bar{\eta}_{\epsilon}\in D^{K}_{\epsilon,\rho}$ holds, if $\epsilon$ is small enough. By (55) and (47)
@@
(57) $\Phi(\eta_{i})=\Phi(\xi_{0})+\frac{1}{2}\Phi^{\prime\prime}(\xi_{0})d(\xi_{0},\eta_{i})^{2}+o(\epsilon)=\Phi(\xi_{0})+\epsilon\Phi^{\prime\prime}(\xi_{0})|e_{i}|^{2}+o(\epsilon)$
@@
where $\gamma:=\min\left\{\int_{{\mathbb{R}}^{n}}U^{p-1}(z)e^{\langle b_{ij},z\rangle}dz|b_{ij}\in{\mathbb{R}}^{n},|b_{ij}|=1\right\}>0.$